ULTRAHIGH-RESOLUTION SPECTROSCOPY AND THE ZEEMAN EFFECT OF THE B^{2}E′ ← X^{2}A′_{2} TRANSITION OF NITRATE RADICAL NO_{3}

\underline{SHUNJI KASAHARA}, KOHEI TADA, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan; MASAAKI BABA, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Hiroshima 731-3194, Japan; EIZI HIROTA, The Graduate University for Advanced Studies, Kanagawa 240-0193, Japan.

The nitrate radical NO_{3} has been known as an important intermediate in chemical reaction in the night atmosphere. The B ^{2}E′ ← X ^{2}A′_{2} transition has been known as an intense absorption in the visible region (T_{0} ≈ 15108 cm^{−1}). The rotational structure of the 0-0 band of the B ^{2}E′ ← X ^{2}A′_{2} transition has been reported by Carter et al.^{a}, but the rotational assignment is still remained because the spectrum is too complicated. In this work, we measured the rotationally resolved high-resolution fluorescence excitation spectra of the 0-0 band of NO_{3} B ^{2}E′ ← X ^{2}A′_{2} transition by crossing a single-mode laser beam perpendicular to a collimated molecular beam in the range of 15080-15135 cm^{−1}. In our spectrum, the typical linewidth was 25 MHz and the absolute wavenumber was calibrated with accuracy 0.0001 cm^{−1} by measurement of the Doppler-free saturation spectrum of iodine molecule and fringe pattern of the stabilized etalon. Additionally, we have observed the change of the spectra with magnetic field. We are trying to assign the observed rotational lines by using the observed Zeeman splitting and the conbination difference in the ground state calculated from the molecular constants reported by Kawaguchi et al. ^{b}

aR. T. Carter, K. F. Schmidt, H. Bitto, and J. R. Huber, Chem. Phys. Lett., 257, 297 (1996).

bK. Kawaguchi, E. Hirota, T. Ishiwata, and I. Tanaka, J. Chem. Phys., 93, 951 (1990).